OpenStudy (anonymous):

20x-2=18x-4 solve for x.

OpenStudy (anonymous):

20x-18x=-4+2 2x=-2 x=-1

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

20x - 18x = -4 + 2 2x = -2 Divide both sides by 2. x = -1 :)